# ♪  Something Just Like TRuST ♪ thanks: An allusion to Coldplay’s “Something Just Like This.” : Toxicity Recognition of Span and Target

### v1

@@ 5 Conclusion
To further research into identifying and mitigating toxicity in LLMs, we have presented TRuST, a dataset of over 298K examples, where about 38% are toxic, and have been labeled for 24 fine-grained target social groups (including "no target"), and toxic spans. A subset of over 11K examples were manually annotated by a diverse group of annotators whose inter-annotator agreements on the three labeling tasks were higher than reported in previous work. We benchmarked 13 methods for detection accuracy and F1 on the three aspects of toxicity. Unsurprisingly, the PLM classifiers outperformed the LLMs. Neither CoT nor reasoning models led to higher LLM performance, thus current methods to improve LLM reasoning do not address social reasoning well, if at all.
The 287K examples that were not human-annotated have been automatically annotated using the best performing models: a RoBERTa PLM for toxicity, a BERT PLM for target group, and SpanBERT for toxic span. This dataset will be released if the paper is accepted for publication. We offer TRuST as a contribution toward more robust methods for toxicity detection, and perhaps to support development of corpora and methods to mitigate toxicity, such as unlearning Chen and Yang (2023); Liu et al. (2024), which to our knowledge has not yet been applied to toxicity.
@@ 6 Limitations
The work presented here carries out only a preliminary investigation of baseline methods for automatic identification of toxicity, target social group and toxic span detection. The LLMs methods did not explore sophisticated prompt engineering. Although the size of the dataset is competitive, it is not sufficiently large to have separate annotations for some important subgroups. Although we attempted to recruit a pool of annotators that was socially diverse, this was limited due to lack of funds to recruit more than six annotators.
@@ Appendix B Experimental Details
We use AdamW optimizer Kingma and Ba (2017) to fine-tune PLMs. We use a batch size of 8. We tried 0.01, 0.0001, 0.00001, 0.05,0.005,0.0005, 0.00005 for learning rate and we chose the best one for each task, we chose the best one based on the validation performance. The best learning rate is 1e-05 for all tasks and embedders. We ran the experiments on a single NVIDIA RTX A6000 GPU and it took 2 hours for each experiment.
@@ Appendix C Annotation Instructions
Creating annotated data for toxic language, where, very generally, a sentence is toxic if it has negative stereotyping, hate speech, racism, psychological threat, sexual harassment, abusive language, sexism, discrimination based on sexual orientation, or any other type of explicit language that might hurt or affect a member of some sociodemographic group badly.
@@
Full Set of Target Groups We have more specific targets within higher level categories. For each category, except individual and disability, we also have other. You should use the label “other $<higher\_level\_target>$” in the case that the group belongs to $<higher\_level\_target>$ and not among the ones listed below (e.g. for the statement “Turkish people are useless”, the target is ethnicity, but there is no Turkish category so it should be other ethnicity)
@@ Appendix F LLM Prompt for Target Group
- no target - other country - united states (notes: must be related to the United States as a whole) - black - asian (note: ethnicity, excludes Chinese) - white - arab - latino (note: ethnicity, includes Hispanic and excludes Mexican) - native american - chinese - mexican - other ethnicity - jewish - muslim - other religion - politics (note: must reference a political figure or party) - woman - man - lgbtq+ - other gender - disability - middle east - refugee (note: includes immigrants) - other For each text, at the end, respond in the following format: Targeted Group: [Group Name] User Prompt: Text: <text>
@@ Appendix G LLM Prompt for Higher Level Target Group
You are an expert in linguistic and sociological analysis. Your task is to determine which social group is targeted by the sentiment or content of each text. For each text, use the provided high level social groups and always select the group that is most directly addressed or implied. If an individual not involved in politics is mentioned, or if no group is targeted, respond with ’Targeted Group: no target’. If multiple groups are targeted with no clear dominance, or if a group is certainly targeted but not clearly identified, respond with ’Targeted Group: other.’ Here are the social groups: - no target - country - ethnicity - religion - politics - gender (note: includes lgbtq+) - disability - other (note: includes politics that references a political figure or party) For each text, at the end, respond in the following format: Targeted Group: [Group Name]
@@ Appendix I Effect of Target Group on Toxicity Detection
Appendix J Accuracy Per Higher Target Group
@@
Appendix K Target Group Accuracy Per Target Group

### v2

@@ 5 Conclusion
The TRuST dataset presented here fills a significant gap in measurement of LLM toxicity by standardizing annotation of toxicity in a manner that allowed us to merge and re-annotate five previous datasets. At 298k examples, the resulting dataset is larger than previous ones. The human-annotated subset of 11k examples has higher inter-annotator reliability on the three labeling tasks than reported in previous work. The quality of the remaining synthetic labels from our highest performing prediction models is close to human quality. Further, PLM classifiers trained on TRuST perform well on an independent toxicity dataset from OpenAI.
Benchmarking of 13 methods using PLMs and diverse LLMs, including in-context learning, automatic prompt optimization, CoT and reasoning models, showed that toxicity prediction of all three annotation elements in TRuST (toxicity, target group, span) is challenging for current LLMs, which underperform PLMs. We suspect that existing reasoning methods for LLMs fail to accommodate the subtle social reasoning involved in toxicity prediction. Through its improved reliability, TRuST should foster progress on toxicity analysis and mitigation methods, such as unlearning chen2023unlearnwantforgetefficient; liu2024rethinkingmachineunlearninglarge, which to our knowledge has not yet been applied to toxicity.
@@ 6 Limitations
The work presented here carries out only a preliminary investigation of baseline methods for automatic identification of toxicity, target social group and toxic span detection. The LLMs methods did not explore sophisticated prompt engineering. Although the size of the dataset is competitive, it is not sufficiently large to have separate annotations for some important subgroups. Although we attempted to recruit a pool of annotators that was socially diverse, this was limited due to lack of funds to pay more than six annotators.
7 Ethical considerations
Because of the nature of our data, there are offensive words in our dataset. All annotators, including volunteers, were aware of this, and they agreed to work on the data. We also include a disclaimer about this on the first page of the paper. Last but not least, when the dataset is public, we will include a warning.
@@ Appendix B Experimental Details
We use AdamW optimizer kingma2017adammethodstochasticoptimization to fine-tune PLMs. We use a batch size of 8. We tried 0.01, 0.0001, 0.00001, 0.05,0.005,0.0005, 0.00005 for learning rate and we chose the best one for each task, we chose the best one based on the validation performance. The best learning rate is 1e-05 for all tasks and embedders. We ran the experiments on a single NVIDIA RTX A6000 GPU and it took 2 hours for each experiment.
@@ Appendix C Annotation Instructions
Creating annotated data for toxic language, where, very generally, a sentence is toxic if it has negative stereotyping, hate speech, racism, psychological threat, sexual harassment, abusive language, sexism, discrimination based on sexual orientation, or any other type of language that might hurt or affect a member of some sociodemographic group badly.
@@
Full Set of Target Groups We have more specific targets within higher level categories. For each category, except individual and disability, we also have other. You should use the label “other <higher_level_target><higher\_level\_target>” in the case that the group belongs to <higher_level_target><higher\_level\_target> and not among the ones listed below (e.g. for the statement “Turkish people are useless”, the target is ethnicity, but there is no Turkish category so it should be other ethnicity)
@@ Appendix F LLM Prompt for Target Group
- no target
- other country
- united states (notes: must be related to the United States as a whole)
- black
- asian (note: ethnicity, excludes Chinese)
- white
- arab
- latino (note: ethnicity, includes Hispanic and excludes Mexican)
- native american
- chinese
- mexican
- other ethnicity
- jewish
- muslim
- other religion
- politics (note: must reference a political figure or party)
- woman
- man
- lgbtq+
- other gender
- disability
- middle east
- refugee (note: includes immigrants)
- other For each text, at the end, respond in the following format: Targeted Group: [Group Name] User Prompt: Text: <text>
@@ Appendix G LLM Prompt for Higher Level Target Group
You are an expert in linguistic and sociological analysis. Your task is to determine which social group is targeted by the sentiment or content of each text. For each text, use the provided high level social groups and always select the group that is most directly addressed or implied. If an individual not involved in politics is mentioned, or if no group is targeted, respond with ’Targeted Group: no target’. If multiple groups are targeted with no clear dominance, or if a group is certainly targeted but not clearly identified, respond with ’Targeted Group: other.’ Here are the social groups: - no target
- country
- ethnicity
- religion
- politics
- gender (note: includes lgbtq+)
- disability
- other (note: includes politics that references a political figure or party)
For each text, at the end, respond in the following format: Targeted Group: [Group Name]
@@ Appendix I Effect of Target Group on Toxicity Detection
Appendix J Span Detection For Sentences vs Subsentence Spans
Appendix K Accuracy Per Higher Target Group
@@
Appendix L Target Group Accuracy Per Target Group
@@
Appendix M Toxicity Detection Accuracy Per Target Group
Figure 6: Toxicity detection accuracy by target group